\begin{thebibliography}{22}
\providecommand{\natexlab}[1]{#1}
\providecommand{\url}[1]{\texttt{#1}}
\expandafter\ifx\csname urlstyle\endcsname\relax
  \providecommand{\doi}[1]{doi: #1}\else
  \providecommand{\doi}{doi: \begingroup \urlstyle{rm}\Url}\fi

\bibitem[Bull(2011)]{bull2011convergence}
Adam~D Bull.
\newblock Convergence rates of efficient global optimization algorithms.
\newblock \emph{Journal of Machine Learning Research}, 12\penalty0 (10), 2011.

\bibitem[Calandriello et~al.(2019)Calandriello, Carratino, Lazaric, Valko, and
  Rosasco]{Calandriello2019Adaptive}
Daniele Calandriello, Luigi Carratino, Alessandro Lazaric, Michal Valko, and
  Lorenzo Rosasco.
\newblock Gaussian process optimization with adaptive sketching: Scalable and
  no regret.
\newblock In \emph{Conference on Learning Theory}, 2019.

\bibitem[Camilleri et~al.(2021)Camilleri, Jamieson, and
  Katz-Samuels]{camilleri2021high}
Romain Camilleri, Kevin Jamieson, and Julian Katz-Samuels.
\newblock High-dimensional experimental design and kernel bandits.
\newblock In \emph{International Conference on Machine Learning}, pages
  1227--1237. PMLR, 2021.

\bibitem[Chen et~al.(2018)Chen, Huang, Wang, Antonoglou, Schrittwieser, Silver,
  and de~Freitas]{chen2018bayesian}
Yutian Chen, Aja Huang, Ziyu Wang, Ioannis Antonoglou, Julian Schrittwieser,
  David Silver, and Nando de~Freitas.
\newblock Bayesian optimization in {A}lpha{G}o.
\newblock \emph{arXiv preprint arXiv:1812.06855}, 2018.

\bibitem[Chowdhury and Gopalan(2017)]{Chowdhury2017bandit}
Sayak~Ray Chowdhury and Aditya Gopalan.
\newblock On kernelized multi-armed bandits.
\newblock In \emph{International Conference on Machine Learning}, pages
  844--853, 2017.

\bibitem[Garivier et~al.(2016)Garivier, Lattimore, and
  Kaufmann]{garivier2016explore}
Aur{\'e}lien Garivier, Tor Lattimore, and Emilie Kaufmann.
\newblock On explore-then-commit strategies.
\newblock \emph{Advances in Neural Information Processing Systems}, 29, 2016.

\bibitem[Gupta et~al.(2022)Gupta, Rana, Venkatesh, et~al.]{gupta2022regret}
Sunil Gupta, Santu Rana, Svetha Venkatesh, et~al.
\newblock Regret bounds for expected improvement algorithms in gaussian process
  bandit optimization.
\newblock In \emph{International Conference on Artificial Intelligence and
  Statistics}, pages 8715--8737. PMLR, 2022.

\bibitem[Jacot et~al.(2018)Jacot, Gabriel, and Hongler]{jacot2018neural}
Arthur Jacot, Franck Gabriel, and Cl{\'e}ment Hongler.
\newblock Neural tangent kernel: Convergence and generalization in neural
  networks.
\newblock \emph{arXiv preprint arXiv:1806.07572}, 2018.

\bibitem[Kassraie and Krause(2022)]{kassraie2022neural}
Parnian Kassraie and Andreas Krause.
\newblock Neural contextual bandits without regret.
\newblock In \emph{International Conference on Artificial Intelligence and
  Statistics}, 2022.

\bibitem[Li and Scarlett(2022)]{li2021gaussian}
Zihan Li and Jonathan Scarlett.
\newblock Gaussian process bandit optimization with few batches.
\newblock In \emph{International Conference on Artificial Intelligence and
  Statistics}, 2022.

\bibitem[Lyu et~al.(2019)Lyu, Yuan, and Tsang]{lyu2019efficient}
Yueming Lyu, Yuan Yuan, and Ivor~W Tsang.
\newblock Efficient batch black-box optimization with deterministic regret
  bounds.
\newblock \emph{arXiv preprint arXiv:1905.10041}, 2019.

\bibitem[Salgia et~al.(2021)Salgia, Vakili, and Zhao]{salgia2021domain}
Sudeep Salgia, Sattar Vakili, and Qing Zhao.
\newblock A domain-shrinking based {B}ayesian optimization algorithm with
  order-optimal regret performance.
\newblock \emph{Conference on Neural Information Processing Systems}, 34, 2021.

\bibitem[Scarlett et~al.(2017)Scarlett, Bogunovic, and
  Cevher]{Scarlett2017Lower}
Jonathan Scarlett, Ilija Bogunovic, and Volkan Cevher.
\newblock Lower bounds on regret for noisy {G}aussian process bandit
  optimization.
\newblock In \emph{Conference on Learning Theory}, pages 1723--1742, 2017.

\bibitem[{Shahriari} et~al.(2016){Shahriari}, {Swersky}, {Wang}, {Adams}, and
  {de Freitas}]{Shahriari2016outofloop}
B.~{Shahriari}, K.~{Swersky}, Z.~{Wang}, R.~P. {Adams}, and N.~{de Freitas}.
\newblock Taking the human out of the loop: A review of {B}ayesian
  optimization.
\newblock \emph{Proceedings of the IEEE}, 104\penalty0 (1):\penalty0 148--175,
  2016.

\bibitem[Snoek et~al.(2012)Snoek, Larochelle, and Adams]{snoek2012practical}
Jasper Snoek, Hugo Larochelle, and Ryan~P Adams.
\newblock Practical bayesian optimization of machine learning algorithms.
\newblock \emph{Advances in neural information processing systems}, 25, 2012.

\bibitem[Srinivas et~al.(2010)Srinivas, Krause, Kakade, and
  Seeger]{srinivas2010gaussian}
Niranjan Srinivas, Andreas Krause, Sham Kakade, and Matthias Seeger.
\newblock Gaussian process optimization in the bandit setting: no regret and
  experimental design.
\newblock In \emph{International Conference on Machine Learning}, pages
  1015--1022. Omnipress, 2010.

\bibitem[Teckentrup(2018)]{Teckentrup2019}
Aretha~L. Teckentrup.
\newblock Convergence of gaussian process regression with estimated
  hyper-parameters and applications in bayesian inverse problems.
\newblock \emph{Available at Arxiv.}, 2018.

\bibitem[Vakili et~al.(2021{\natexlab{a}})Vakili, Bromberg, Garcia, Shiu, and
  Bernacchia]{vakili2021uniform}
Sattar Vakili, Michael Bromberg, Jezabel Garcia, Da-shan Shiu, and Alberto
  Bernacchia.
\newblock Uniform generalization bounds for overparameterized neural networks.
\newblock \emph{arXiv preprint arXiv:2109.06099}, 2021{\natexlab{a}}.

\bibitem[Vakili et~al.(2021{\natexlab{b}})Vakili, Khezeli, and
  Picheny]{vakili2020information}
Sattar Vakili, Kia Khezeli, and Victor Picheny.
\newblock On information gain and regret bounds in {Gaussian} process bandits.
\newblock In \emph{International Conference on Artificial Intelligence and
  Statistics}, pages 82--90, 2021{\natexlab{b}}.

\bibitem[Vakili et~al.(2021{\natexlab{c}})Vakili, Scarlett, and
  Javidi]{vakili2021open}
Sattar Vakili, Jonathan Scarlett, and Tara Javidi.
\newblock Open problem: Tight online confidence intervals for {RKHS} elements.
\newblock In \emph{Conference on Learning Theory}, pages 4647--4652. PMLR,
  2021{\natexlab{c}}.

\bibitem[Vakili et~al.(2022)Vakili, Scarlett, Shiu, and
  Bernacchia]{vakili2022improved}
Sattar Vakili, Jonathan Scarlett, Da-shan Shiu, and Alberto Bernacchia.
\newblock Improved convergence rates for sparse approximation methods in
  kernel-based learning.
\newblock In \emph{International Conference on Machine Learning}, 2022.

\bibitem[Valko et~al.(2013)Valko, Korda, Munos, Flaounas, and
  Cristianini]{Valko2013kernelbandit}
Michal Valko, Nathan Korda, R\'{e}mi Munos, Ilias Flaounas, and Nello
  Cristianini.
\newblock Finite-time analysis of kernelised contextual bandits.
\newblock In \emph{Conference on Uncertainty in Artificial Intelligence}, page
  654–663, 2013.

\end{thebibliography}

\documentclass[final,12pt]{colt2022} 

\title[Noise-Free Kernel-Based Bandit]{Open Problem: Regret Bounds for Noise-Free Kernel-Based Bandits}
\usepackage{times}

\medskip
\coltauthor{%
 \Name{Sattar Vakili} \Email{sattar.vakili@mtkresearch.com}\\
 \addr MediaTek Research
}

\def\Oc{\mathcal{O}}
\def\Oct{\tilde{\mathcal{O}}}

\usepackage{setspace}

\usepackage[utf8]{inputenc} 
\usepackage[T1]{fontenc}    
\usepackage{hyperref}  
\hypersetup{
    colorlinks=true,
    linkcolor=blue,
    citecolor =blue,
    filecolor=magenta,      
    urlcolor=magenta,
}
\usepackage{url}            
\usepackage{booktabs}       
\usepackage{amsfonts}       
\usepackage{nicefrac}       
\usepackage{microtype}      
\usepackage{lipsum}
\usepackage{graphicx}

\usepackage{amsmath}
\usepackage{amsfonts}
\usepackage{algorithm}
\usepackage{algorithmic}
\usepackage{bm}
\usepackage{amssymb}
\usepackage{color}
\usepackage{xcolor}
\usepackage{hyperref}
\usepackage{amsmath, amssymb, graphicx, url}

\usepackage{cleveref}
\crefformat{section}{\S#2#1#3} 
\crefformat{subsection}{\S#2#1#3}
\crefformat{subsubsection}{\S#2#1#3}

\usepackage{mathrsfs}
\usepackage{makecell}

\def\Ib{\bm{I}}

\def\Hc{\mathcal{H}}

\def\TP{\top}

\def\argmax{\text{argmax}}

\def\Kb{\mathbf{K}}
\def\kb{\mathbf{k}}

\def\yb{\mathbf{y}}
\def\Ib{\mathbf{I}}

\def\Rr{\mathbb{R}}

\def\E{\mathbb{E}}

\def\Xc{\mathcal{X}}
\def\Rc{\mathcal{R}}
\def\Oc{\mathcal{O}}
\def\Oct{\tilde{\mathcal{O}}}

\newtheorem{assumption}{Assumption}

\begin{document}

\maketitle

\begin{abstract}%
  Kernel-based bandit is an extensively studied black-box optimization problem, in which the objective function is assumed to live in a known reproducing kernel Hilbert space. While nearly optimal regret bounds (up to logarithmic factors) are established in the noisy setting, surprisingly, less is known about the noise-free setting (when the exact values of the underlying function is accessible without observation noise). We discuss several upper bounds on regret; none of which seem order optimal, and provide a conjecture on the order optimal regret bound. 

\end{abstract}

\begin{keywords}%
Kernel-based bandit, Bayesian optimization, Gaussian processes
\end{keywords}

\section{Introduction}

Black-box optimization of a (possibly non-convex) function from expensive evaluations is a ubiquitous problem in machine learning, including both academic research and industrial applications. Those include A/B testing, hyperparameter tuning \citep[including AlphaGo,][]{chen2018bayesian}, robotics, environmental monitoring, and more \citep{Shahriari2016outofloop}. A particularly successful approach is based on the use of Gaussian process modeling, due to its versatility, modeling power, and ability to provide uncertainty estimates. A mathematical formulation of the problem under the bandit setting has been extensively studied in the literature. Nearly optimal regret bounds are established in the noisy setting. Surprisingly, the problem of order optimal regret bounds remains open under the noise-free setting. Here, we overview the existing results and give a formal description of the open problem, aspiring to motivate solutions.

\section{Problem Setup}\label{sec:setup}

Consider the sequential optimization of an objective function $f:\Xc\rightarrow\Rr$, where $\Xc\subset\Rr^d$ is a compact set. A learning algorithm is allowed to collect a sequence of observations $\{(x_i, y_i)\}_{i=1}^\infty$, where $y_i=f(x_i)$ in the noise-free setting, and $y_i=f(x_i)+\epsilon_i$ in the noisy setting with $\epsilon_i$ being a well behaved observation noise. The objective is to get as close as possible to the maximum of~$f$. The performance of the algorithm is measured in terms of (cumulative) regret, defined as the cumulative loss in the values of the objective function at observation points, compared to a global maximum:
\begin{eqnarray}
\Rc(N) = \sum_{i=1}^N \left(f(x^*) - f(x_i)\right), 
\end{eqnarray}
where $x^*\in \argmax_{x\in \Xc} f(x)$ is a global maximum.  The regularity assumption on $f$ which makes the problem tractable is given next. 

\paragraph{The RKHS and the regularity assumption on $f$:} Consider a positive definite kernel $k:\Xc \times \Xc\rightarrow \Rr$ with respect to a finite Borel measure. Let $\Hc_k$ denote the reproducing kernel Hilbert space (RKHS) corresponding to $k$, defined as a Hilbert space equipped with an inner product $\langle.,.\rangle_{\Hc_k}$ satisfying the following: $k(.,x)\in \Hc_k$, $\forall x\in \Xc$, and $\langle f,k(.,x)\rangle_{\Hc_k}=f(x)$, $\forall x\in\Xc, \forall f \in \Hc_k$ (reproducing property). The typical assumption in kernel-based models is that the objective function~$f$ satisfies $f\in\Hc_k$ for a known kernel $k$. 
\begin{assumption}\label{ass_norm}
Assume $f$ is fixed with $\|f\|_{\Hc_k}\le C_k$, for some $C_k>0$, where $\|f\|_{\Hc_k}=\sqrt{\langle f,f\rangle_{\Hc_k}}$. 
\end{assumption}


Under Assumption~\ref{ass_norm}, the sequential optimization problem given above is often referred to as that of kernel-based (kernelized) bandit, Gaussian process (GP) bandit, or Bayesian optimization. The latter two terms are motivated by the algorithm design which often employs a GP surrogate model. 

\paragraph{GP surrogate  model:}

It is useful for algorithm design to employ a zero-mean GP model $F$ with kernel $k$, which provides a surrogate posterior mean (prediction) and a surrogate posterior variance (uncertainty estimate) for the kernel-based model.  Defining $\mu_n(x) = \E\big[F(x)|\{(x_i,y_i)\}_{i=1}^{n}\big]$ and $\sigma_n^2(x) = \E\big[(F(x) - \mu_n(x))^2|\{(x_i,y_i)\}_{i=1}^{n}\big]$, it is well known that
$
\mu_n(x) =  \kb_n^{\TP}(x)  (\lambda^2 \Ib_n+\Kb_n)^{-1}\yb_n$ and 
$\sigma_n^2(x) = k(x,x) - \kb_n^{\TP}(x)  (\lambda^2 \Ib_n+\Kb_n)^{-1}\kb_n(x), 
$
where
$\yb_n=[y_1,\dots, y_n]^{\top}$,
$\kb_n(x) = [k(x,x_1), \dots, k(x,x_n)]^{\TP}$, $\Kb_n$ is the positive definite kernel matrix $[\Kb_n]_{i,j} = k(x_i, x_j)$
, $\Ib_n$ is the identity matrix of dimension $n$, and $\lambda^2$ is the variance of a zero mean Gaussian surrogate distribution for the noise. In the noise-free setting, one may set $\lambda=0$ in the above expressions. 

The Mat\'ern family of kernels is perhaps the most commonly used in practice~\citep[see, e.g.,][]{snoek2012practical, Shahriari2016outofloop} and theoretically interesting~\citep[see, e.g.,][]{ srinivas2010gaussian,bull2011convergence} family of kernels. It is known that the RKHS corresponding to a Mat\'ern kernel is equivalent to a Sobolev space~\citep[see, e.g.,][]{Teckentrup2019} that provides an intuitive interpretation of the function class based on its smoothness. Furthermore, the Mat{\'e}rn kernels may be insightful for the theory of neural networks, as it has been shown that the neural tangent kernel~\citep{jacot2018neural} is equivalent to a Mat{\'e}rn kernel~\citep{vakili2021uniform}. Following the literature, we also emphasize the Mat{\'e}rn family in the formulation of the open problem.

\section{Overview of the Results in the Noisy Setting}


We first overview the main results for the kernel-based bandit problem in the noisy setting and under Assumption~\ref{ass_norm}. 
Classical algorithms such as GP upper confidence bound \citep[GP-UCB,][]{srinivas2010gaussian, Chowdhury2017bandit}, Thompson sampling~\citep[GP-TS,][]{Chowdhury2017bandit}, and expected improvement~\citep[EI,][]{gupta2022regret} sequentially select the observation points based on a score referred to as {acquisition function}. The best known regret bounds for these algorithms
scales as $\Oct(\Gamma_{k,\lambda}(N)\sqrt{N})$ over $N$ steps (see the references above), where $\Gamma_{k,\lambda}(N)$ is a kernel specific complexity term, 
referred to as the \emph{maximal information gain} between the noisy observations and the latent GP surrogate model~\citep{srinivas2010gaussian}:
\begin{eqnarray}
\Gamma_{k,\lambda}(n) =\sup_{\{x_i\}_{i=1}^n\subset\Xc}\frac{1}{2}\log\det(\Ib_n+\frac{1}{\lambda^2}\Kb_n).
\end{eqnarray}
That is also nearly equivalent to the \emph{effective dimension} of the kernel for a dataset of finite size $n$~\citep{Calandriello2019Adaptive}. Kernel specific bounds on $\Gamma_{k,\lambda}(n)$, with a focus on Mat{\'e}rn family, are provided in~\cite{srinivas2010gaussian, vakili2020information, vakili2021uniform, kassraie2022neural}. 

The $\Oct(\Gamma_{k,\lambda}(N)\sqrt{N})$\footnote{The notations $\Oc$ and $\Oct$ are used for mathematical order, and that up to hiding logarithmic factors, respectively.} scaling is not tight in general, and may even fail to be sublinear in many cases of interest, since $\Gamma_{k,\lambda}(N)$ may grow faster than $\sqrt{N}$. It remains an open problem whether the suboptimal regret bounds of these acquisition based algorithms is a fundamental limitation or a shortcoming of their proof~\citep[see][for the details]{vakili2021open}.

On discrete domains, the SupKernelUCB algorithm was shown to have an $\Oct(\sqrt{\Gamma_{k,\lambda}(N)N})$ regret bound~\citep{Valko2013kernelbandit}.
SupKernelUCB is not considered to be practical. Several more practical algorithms (which also apply to continuous domains) with $\Oct(\sqrt{\Gamma_{k,\lambda}(N)N})$ regret have been proposed recently: a tree-based domain-shrinking algorithm~\citep[GP-ThreDS,][]{salgia2021domain}, Robust Inverse Propensity Score for experimental design~\citep[RIPS,][]{camilleri2021high}, batched pure exploration~\cite[BPE,][]{li2021gaussian}, and its sparse version~\citep[S-BPE,][]{vakili2022improved}.

Substituting the bounds on $\Gamma_{k,\lambda}(N)$ for the Mat\'ern kernels leads to $\Rc(N)=\Oct(N^{\frac{\nu+d}{2\nu+d}})$, where $\nu$ is the smoothness parameter of the kernel. \cite{Scarlett2017Lower} proved a matching (up to logarithmic factors) $\Omega(N^{\frac{\nu+d}{2\nu+d}})$ lower bound on the regret, which shows the order optimality of the $\Oct(\sqrt{\Gamma_{k,\lambda}(N)N})$ regret bounds in this case. 


\section{Regret Bounds in the Noise-Free Setting}

As outlined above, there are many results on the kernel-based bandit problem under the noisy setting, several of which achieving nearly optimal regret bounds. Under the noise-free setting,
\cite{lyu2019efficient} is the only work providing bounds on the cumulative regret, while \cite{bull2011convergence} provided nearly optimal bounds on the simple regret. In this section, we first overview these two works, and then provide a formal description of the open problem of cumulative regret under the noise-free setting. 

\subsection{From Noisy to Noise-Free}

One straightforward way to establish regret bounds under the noise-free setting is to set the noise equal to zero in the results under the noisy setting. 
\cite{lyu2019efficient} took this approach and showed an
$\Oc(\sqrt{\Gamma_{k,\lambda}(N) N})$ bound on the cumulative regret for GP-UCB. 
The analysis mainly follows that of GP-UCB under the noisy setting~\citep{srinivas2010gaussian}, except for using tighter confidence intervals for GP models under the noise-free setting. In the case of Mat{\'e}rn kernels, this implies an $\Rc(N)=\Oct(N^{\frac{\nu+d}{2\nu+d}})$ regret that always grows faster than $\sqrt{N}$, even for highly smooth kernels (large $\nu$). As we will see next, this regret bound is not order optimal in the noise-free setting.

\subsection{Explore then Commit}

Consider the simple regret variation of the kernel-based bandit problem. In this variation, after $n$ (possibly adaptive) observations $\{(x_i,y_i)\}_{i=1}^{n}$, the algorithm selects a candidate maximizer $\hat{x}_n$, and its performance is measured in terms of simple regret defined as $r_n=f(x^*)-f(\hat{x}_n)$. 
Focusing on Mat\'ern family and simple regret, \cite[][]{bull2011convergence} proved an $r_n=\Oct(n^{-(\min\{\nu,1\})/d})$ regret bound for the EI algorithm (see their Theorem $2$), that translates to an $R(N)=\Oc(N^{(d-1)/d})$ bound on its cumulative regret when $\nu\ge 1$ (which is typically the case, except for the Laplace kernel where $\nu=\frac{1}{2}$).
Furthermore, they showed that EI mixed with pure exploration achieves
an $r_n=\Oct(n^{-\frac{\nu}{d}})$ simple regret (see their Theorem $5$), which cannot be improved (see their Theorem~$1$). 
The cumulative regret $R(N)$ for this algorithm, however, grows linearly in $N$, as a consequence of pure exploration. The question of cumulative regret appears more challenging due to the
exploration-exploitation tradeoff, which is inherent to the bandit problems.

Inspired by the results of \cite{bull2011convergence}, and using the \emph{explore then commit} technique~\citep[e.g., see][]{garivier2016explore}, we can design a simple algorithm with sublinear regret. Specifically, consider an algorithm which performs pure exploration by, e.g., choosing nearly uniform points across the domain, up to step $N_0$. The algorithm then commits to the best observation point based on the GP prediction: $\hat{x}_{N_0}=\arg\max_{x\in\Xc}\mu_{N_0}(x)$ (with $\lambda=0$ in the expression of $\mu_{N_0}$), for the remaining of the steps: $x_i=\hat{x}_{N_0}, \forall i>N_0$ . Choosing the optimum value for $N_0$, leads to a cumulative regret of $\Rc(N)= \Oc(N^{\frac{d}{\nu+d}})$. Note that this regret bound is tighter than the one in~\cite{lyu2019efficient} given above, indicating that the analytical techniques used in the noisy setting may not be suitable for the noise-free setting.


\subsection{Open Problem}

\paragraph{Problem~1}
Consider the kernel-based bandit problem given in Section~\ref{sec:setup} under the noise-free setting ($\epsilon_i=0, \forall i$). What is the lowest growth rate of $R(N)$ with $N$? 
As a specific case of interest, when the kernel $k$ is a Mat{\'e}rn kernel with smoothness parameter $\nu$ and $R(N)=\Oct(N^{\alpha})$, what is the smallest value of $\alpha$, achievable by a learning algorithm?

\subsection{Discussion}

In the case of Mat{\'e}rn kernels, as outlined above, the existing upper bounds include $\Oct(N^{\frac{\nu+d}{2\nu+d}})$~\citep{lyu2019efficient}, $\Oc(N^{\frac{d-1}{d}})$~\citep[EI,][]{bull2011convergence}, and $\Oct(N^{\frac{d}{\nu+d}})$ (explore then commit). We however conjecture that the following regret bound is achievable:
\begin{eqnarray}
\Rc(N)= 
\begin{cases}
    \Oc(N^{(d-\nu)/d}),~~~ &\text{when}~d>\nu,\\
    \Oc\left(\log (N)\right), &\text{when}~d=\nu,\\
    \Oc(1), &\text{when}~d<\nu.
\end{cases}
\end{eqnarray}


We here give an informal reasoning for this claim. Following standard UCB-based bandit algorithm techniques, it can be shown that the GP-UCB
algorithm attains $\Rc(N)=\Oc(C_k\sum_{i=1}^N\sigma_{i-1}(x_i))$. Let $\{x^*_i\}_{i=1}^n\in\arg\max_{\{x_i\}_{i=1}^n\subset \Xc}\sum_{i=1}^n\sigma_{i-1}(x_i)$ and $\Theta^*_n= \sum_{i=1}^n\sigma_{i-1}(x^*_i)$. We then have $\Rc(N)=\Oc(C_k\Theta^*_N)$, which reduces the problem to bounding $\Theta^*_n$. We conjecture that under certain mild regularity of the domain, $\{x^*_i\}_{i=1}^n$ are distributed nearly uniformly across the domain, in which case $\sigma_{i-1}(x^*_i)=\Oc(i^{-\frac{\nu}{d}})$~\cite[see, e.g.,][Lemma~$3.8$]{Teckentrup2019}. That, when summed over $i$, leads to $\Theta^*_n=\Oc(n^{\frac{d-\nu}{d}})$ when $d>\nu$, $\Theta^*_n=\Oc\left(\log(n)\right)$ when $d=\nu$, and $\Theta^*_n=\Oc(1)$ when $d<\nu$.



\bibliography{references}

\appendix



\end{document}